0.4i'+10i=40-sin35t


1
Expert's answer
2021-12-13T13:43:50-0500
"0.4i'+10i=40-\sin35t""i'+25i=100-2.5\sin35t"

Integrating factor


"\mu(t)=e^{\int25dt}=e^{25t}"

"e^{25t}(i'+25i)=100e^{25t}-2.5e^{25t}\sin35t"

"d(e^{25t}i)=(100e^{25t}-2.5e^{25t}\sin35t)dt"

Integrate


"\int d(e^{25t}i) =\int (100e^{25t}-2.5e^{25t}\sin35t)dt"

"I_1=\int e^{25t}\sin35tdt"

"u=\sin35t, du=35\cos35tdt"

"dv=e^{25t}dt, v=\int e^{25t}dt=0.04e^{25t}"

"I_1=\int e^{25t}\sin35tdt"

"=0.04e^{25t}\sin35t-1.4\int e^{25t}\cos35tdt"

"\int e^{25t}\cos35tdt"

"u=\cos35t, du=-35\sin35tdt"

"dv=e^{25t}dt, v=\int e^{25t}dt=0.04e^{25t}"

"\int e^{25t}\cos35tdt=0.04e^{25t}\cos35t""+1.4\int e^{25t}\sin35tdt"

"\int e^{25t}\sin35tdt=0.04e^{25t}\sin35t-"

"-0.056e^{25t}\cos35t-1.96\int e^{25t}\sin35tdt"

"2.96\int e^{25t}\sin35tdt=0.04e^{25t}\sin35t"

"-0.056e^{25t}\cos35t"

"\int e^{25t}\sin35tdt=\dfrac{2}{148}e^{25t}\sin35t-\dfrac{2.8}{148}e^{25t}\cos35t"

"e^{25t}i=4e^{25t}-\dfrac{5}{148}e^{25t}\sin35t+\dfrac{7}{148}e^{25t}\cos35t+C"

"i=4-\dfrac{5}{148}\sin35t+\dfrac{7}{148}\cos35t+Ce^{-25t}"